OpenStudy (sleepyjess):

Eliminate the parameter. \(x = t^2 + 1, y = t^2-1\) \(y = x + 2, x \ge 1\\ y = x - 2, x \ge 1\\ y = x^2 - 2, x \ge 1\\ y = x^2 + 2, x \ge 1\)

OpenStudy (sleepyjess):

I'm getting stuck on this one @phi @xapproachesinfinity

OpenStudy (phi):

if x= t^2 + 1, x is positive (right ?). In fact \( t^2 \ge 0 \) , so \(x \ge 0+1 \) solving for t^2 : t^2= x-1 putting that in the 2nd equation y = t^2 -1 y = x-1 - 1 y= x-2; x \( \ge\)1

OpenStudy (sleepyjess):

Aha! That is what I was doing wrong, I was attempting to solve for t instead of \(t^2\) >.<

OpenStudy (xapproachesinfinity):

eh no just solve fo t^2 lol

OpenStudy (phi):

yes, you must be sly like a fox sometimes

OpenStudy (sleepyjess):

haha thanks :)

OpenStudy (xapproachesinfinity):

hehe true!

OpenStudy (xapproachesinfinity):

don't worry i feel dumb sometimes trying such thing, in fact, it happens to me quite many times lol